. The vapor pressure of ethanol is 400 mmHg at 63.5
Answer #1

21

by using this equation

P2 = 108.5 mmHg = 109 mmHg

answer B is correct

putting values

ln(P2 / 400) = ((39.3*10^3)/8.314 )* (1/(273+63.5) - (1/(273+34.9) )

ln(P2 / 400) = - 1.30

P2 = 400 * e^(-1.30)

P2 = 109.01271721360504 mm Hg

AS 1 TORR = 1 mm Hg ; hence there would be no difference in answer other than units .